### v1

# Mechanistic Fine-tuning for In-context Learning

Abstract
In-context Learning (ICL) utilizes structured demonstration-query inputs to induce few-shot learning on Language Models (LMs), which are not originally pre-trained on ICL-style data. To bridge the gap between ICL and pre-training, some approaches fine-tune LMs on large ICL-style datasets by an end-to-end paradigm with massive computational costs. To reduce such costs, in this paper, we propose Attention Behavior Fine-Tuning (ABFT), utilizing the previous findings on the inner mechanism of ICL, building training objectives on the attention scores instead of the final outputs, to force the attention scores to focus on the correct label tokens presented in the context and mitigate attention scores from the wrong label tokens. Our experiments on 9 modern LMs and 8 datasets empirically find that ABFT outperforms in performance, robustness, unbiasedness, and efficiency, with only around 0.01% data cost compared to the previous methods. Moreover, our subsequent analysis finds that the end-to-end training objective contains the ABFT objective, suggesting the implicit bias of ICL-style data to the emergence of induction heads. Our work demonstrates the possibility of controlling specific module sequences within LMs to improve their behavior, opening up the future application of mechanistic interpretability.
Mechanistic Fine-tuning for In-context Learning
[…]
1 Introduction
In-Context Learning (ICL) (Radford et al., 2019; Dong et al., 2022) is an emerging few-shot learning paradigm where only a concatenation of few-shot demonstrations and a query is needed to conduct the specified task on the query, requiring only feed-forward calculation on the pre-trained Language Models (LMs), as shown in Fig. 1 (A, B). However, trained on natural language data, LMs may face a distribution gap with ICL-style inputs, potentially hindering ICL performance. Therefore, some prior studies (see §2) try to bridge such a gap by fine-tuning LMs on the ICL-style data on end-to-end paradigms, with enormous datasets and calculation cost, preventing practical application, especially on the scaling Large LMs (LLMs).
Therefore, in this paper, we try to propose an efficient fine-tuning approach towards better ICL performance, utilizing some previous observations on the inner mechanisms of ICL. In detail, we focus on the Induction Heads in Transformer-based LMs, which are a set of critical attention heads towards ICL, where the attention scores of the last token in the ICL input (where the predictions are generated) are dominant on the label tokens in the demonstrations, as shown in Fig. 1 (C), for the clue that the tendency of attention scores from induction heads influences the tendency of prediction synchronously Reddy (2024); Cho et al. (2025a) (e.g., if the attention scores of the induction heads focus on the label token “negative” in the context, then the prediction is biased towards “negative”).
Consequently, we can directly control the attention scores to make the induction heads focus on the correct label tokens for correct predictions. Given such an objective, as shown in Fig. 1 (C, D), we propose Attention Behavior Fine-Tuning (ABFT), calculating fine-tuning objective (loss function) only on the attention scores of induction heads, to mitigate “wrong” attention score focusing on the wrong label tokens, and promote “correct” attention score focusing on the correct label tokens. On such an objective, we fine-tune only the and projection matrices of every attention head, with an ICL-style training set of only a few hundred samples, and only a few million of the parameters with gradient activated, which is highly efficient compared to previous works.
Our experiments on 9 modern (L)LMs and 8 downstream datasets demonstrate that ABFT significantly improves ICL performance with satisfactory efficiency, robustness, unbiasedness, and harmlessness, even outperforming previous works of end-to-end fine-tuning the whole model on massive datasets that are approximately larger than ours. Moreover, our analysis finds that the ABFT objective is an implicit bias of direct end-to-end training objective on ICL-style data, suggesting that the causal language modeling on the ICL-style data may naturally evoke the emergence of induction heads.
Our contribution can be summarized as:
We propose Attention Behavior Fine-Tuning (ABFT), which efficiently fine-tunes LMs on ICL inputs using attention-based objectives without supervision on the final output.
Subsequent analysis indicates that the training objective of ABFT is implicitly encompassed by the end-to-end training objective on ICL-style data, suggesting that these data may naturally evoke the induction heads, which enhances the previous works on the inner mechanism of ICL.
Also, we prototypically confirm the possibility of optimizing model performance directly by controlling the intermediate behavior, without any error propagation from the output. This is a hint toward Mechanistic Controllability, a valuable future of mechanistic interpretability.
2 Background & Related Works
In-context Learning.
LM Warm-up for ICL.
Since LMs are typically pre-trained on plain natural language data instead of ICL-style data, it can be expected that a distribution gap between the pre-training and ICL testing occurs to prevent optimal performance. Therefore, some works focus on tuning LMs on the ICL data Chen et al. (2022); Min et al. (2022); Mishra et al. (2022); Wang et al. (2022); Wei et al. (2023). Even effective, these works need gradient-based whole-model and full-precision training on large datasets, making it hard to adapt to real-world applications due to the calculation overhead, and misaligning with the low-resource purpose of ICL.
Induction Heads in ICL Inference.
As shown in Fig. 1 (C), it has been found that some attention heads (called induction heads) in LMs have a nontrivial influence on ICL inference Elhage et al. (2021); Olsson et al. (2022); Singh et al. (2024); Reddy (2024); Cho et al. (2025a), where the attention scores from the last token of the query (the location for prediction, e.g., the last “: ” in Fig. 1 (A), as the attention query) concentrate on the label tokens presented in the demonstrations (e.g., “positive”s and “negative”s in Fig. 1 (A), as the attention key). Attention connections from induction heads transfer label information from the demonstration to the output, biasing predictions toward labels with higher attention scores. Consequently, the accuracy of ICL prediction critically depends on whether these attentions are on the correct labels.
3 Attention Behavior Fine-tuning
Given the inspiration from the previous works, where the ICL predictions are biased towards the more attention-score-concentrated labels in the induction heads, in this paper, as shown in Fig. 1, we propose Attention Behavior Fine-Tuning (ABFT), a novel low-resource fine-tuning method to induce attention scores to focus on the correct labels.
Method Pipeline.
Globally, ABFT utilizes such a pipeline: (1) Dataset Building: from a selected downstream dataset, we build a training set composed of ICL-style sequences as shown in Fig. 1 (A). (2) Feed-forward Calculation: For each training sample, as shown in Fig. 1 (B), we conduct a standard feed-forward calculation on the pre-trained LM, and collect the attention matrices of all the attention heads in all the layers. (3) Loss Calculation: For each attention matrix, we only focus on the attention scores of the last token (i.e., the last row of the attention matrix), where the predictions of queries are made. As shown in Fig. 1 (D), we first filter (detailed below) the non-induction head out, and return a loss of for these heads. Then, for the remaining induction heads, we calculate a loss function composed of a of attention scores on wrong labels and a of attention scores on correct labels (detailed below). (4) Back Propagation: We back-propagate the calculated loss only to the and matrices of every attention head, and update the model parameters.
Induction Head Filter.
As shown in Fig. 1 (C), we skip the attention matrices where the attention scores of the last token do not dominate on the label tokens. To identify the attention matrices to skip, in detail, given an attention matrix , where the is the input token sequence length, as mentioned before, we focus on the last row . Given the position index of label tokens as , we calculate the attention score sum on these label tokens as .111For the case shown in Fig. 1 (C), (0-started), and is the sum of the values of the red-highlighted bar and the blue-highlighted bar. Then, we set a threshold , if , we assert the attention head of score is an induction head, and vice versa. We will verify the necessity and benefits of this induction head filter in §5.2.
Model | Param. | Method | SST2 | MR | FP | SST5 | TREC | SUBJ | TEE | TEH | Average |
GPT2-L | 812M | Vanilla | 56.35 | 61.13 | 42.09 | 29.69 | 35.45 | 49.12 | 39.06 | 54.00 | 45.86 |
MetaICL | 85.94 | 80.96 | 37.30 | 42.09 | 33.98 | 50.49 | 45.41 | 54.20 | 53.80 | ||
PICL | 74.70 | 73.34 | 54.49 | 33.79 | 32.91 | 51.37 | 47.46 | 53.42 | 52.68 | ||
Calibrate | 56.35 | 60.94 | 36.91 | 25.10 | 34.08 | 49.32 | 36.72 | 53.52 | 44.12 | ||
ABFT | 88.181.54 | 85.400.95 | 81.301.69 | 36.843.61 | 50.242.49 | 61.994.39 | 46.513.11 | 55.200.20 | 63.21 | ||
GPT2-XL | 1.61B | Vanilla | 67.87 | 69.53 | 51.07 | 30.66 | 35.25 | 50.98 | 42.68 | 53.03 | 50.13 |
PICL | 74.80 | 74.32 | 51.17 | 32.71 | 33.20 | 51.46 | 47.95 | 53.42 | 52.38 | ||
Calibrate | 68.16 | 75.00 | 36.43 | 28.52 | 35.55 | 50.10 | 39.26 | 51.56 | 48.07 | ||
ABFT | 87.921.47 | 86.521.50 | 87.670.45 | 37.552.67 | 51.832.73 | 75.072.96 | 60.015.38 | 55.350.04 | 67.74 | ||
Falcon3 | 7.46B | Vanilla | 91.11 | 92.77 | 85.35 | 46.00 | 50.00 | 62.60 | 60.55 | 52.05 | 67.55 |
Calibrate | 90.53 | 93.07 | 82.71 | 44.04 | 54.30 | 62.40 | 54.79 | 51.76 | 66.70 | ||
ABFT | 92.140.21 | 92.170.04 | 96.140.36 | 47.320.16 | 75.810.19 | 94.870.82 | 67.970.25 | 70.340.22 | 79.59 | ||
Llama3 | 8.03B | Vanilla | 89.35 | 92.87 | 75.78 | 44.24 | 55.76 | 62.30 | 57.91 | 54.59 | 66.60 |
Calibrate | 90.04 | 93.36 | 43.95 | 41.60 | 54.39 | 65.23 | 54.79 | 52.83 | 62.02 | ||
ABFT | 93.141.39 | 92.500.84 | 94.021.72 | 52.102.31 | 73.091.11 | 92.701.44 | 72.021.81 | 72.025.87 | 80.20 | ||
DeepSeek-R1 Dist. Qwen 4-bit, LoRA | 14.8B | Vanilla | 90.92 | 91.21 | 92.18 | 46.97 | 62.50 | 66.60 | 66.60 | 63.87 | 72.61 |
Calibrate | 90.04 | 91.41 | 92.68 | 46.09 | 61.62 | 65.43 | 65.33 | 62.30 | 71.86 | ||
ABFT | 93.511.22 | 91.850.34 | 91.174.44 | 46.091.27 | 69.142.54 | 92.923.27 | 69.820.00 | 71.193.42 | 78.21 | ||
Qwen2.5 4-bit, LoRA | 32.8B | Vanilla | 93.85 | 94.43 | 86.23 | 47.17 | 58.40 | 87.50 | 65.14 | 69.63 | 75.29 |
Calibrate | 93.75 | 94.82 | 74.22 | 44.82 | 58.79 | 84.08 | 63.96 | 63.96 | 72.30 | ||
ABFT | 94.920.00 | 94.830.10 | 94.040.00 | 48.490.05 | 69.240.10 | 96.000.00 | 69.290.24 | 70.310.00 | 79.64 | ||
SimpleScaling s1.1 4-bit, LoRA | 32.8B | Vanilla | 94.82 | 94.24 | 91.11 | 50.20 | 69.63 | 89.65 | 68.36 | 72.17 | 78.77 |
Calibrate | 94.43 | 93.85 | 88.96 | 48.63 | 68.07 | 89.26 | 68.55 | 72.36 | 78.02 | ||
ABFT | 94.920.10 | 94.290.05 | 96.000.00 | 49.950.05 | 72.460.39 | 95.710.10 | 71.730.05 | 73.100.04 | 81.02 | ||
Llama3 4-bit, LoRA | 43.2B | Vanilla | 93.26 | 94.04 | 73.92 | 49.41 | 58.98 | 71.58 | 62.60 | 66.70 | 71.31 |
Calibrate | 95.02 | 93.07 | 54.20 | 44.53 | 59.08 | 72.56 | 61.03 | 65.82 | 68.16 | ||
ABFT | 95.020.10 | 93.850.10 | 94.870.05 | 48.100.14 | 64.700.14 | 90.090.05 | 69.240.29 | 70.850.15 | 78.34 | ||
Llama3 4-bit, LoRA | 55.6B | Vanilla | 93.94 | 92.19 | 78.81 | 51.37 | 67.19 | 66.70 | 56.44 | 60.94 | 70.95 |
Calibrate | 92.29 | 92.77 | 69.92 | 50.49 | 68.75 | 65.92 | 58.11 | 62.70 | 70.12 | ||
ABFT | 94.530.10 | 93.410.14 | 93.210.44 | 49.020.78 | 71.780.58 | 92.680.48 | 70.760.64 | 70.750.34 | 79.52 |
Loss Function.
As shown in Fig. 1 (D), given an attention matrix , if judged as a non-induction head by the aforementioned head filter, the loss for is assigned to . Else, we conduct the following calculation: given the position index of label tokens consistent with the ground-truth label of the query as , and the others ,222For the case shown in Fig. 1 (C), (the position of “positive”), (the position of “negative”). we calculate the loss from the last row () of as:
That is, as shown in Fig. 1 (D), we the “wrong” attention scores towards the label tokens different from the query’s ground-truth with magnitude , and the “correct” attention scores with magnitude . These two terms in the loss function may seem redundant, but we will demonstrate in §5.3 that they actually contain antagonistic implicit biases, therefore, the factors and should be balanced well.
Why not End-to-end LoRA?
Intuitively, directly adding LoRA bypasses Hu et al. (2022) to the trained projection matrices and fine-tuning them on an end-to-end training objective is also a possible approach. However, in end-to-end LoRA, gradients are propagated from the output logits of LMs, which requires that the final layer of the LMs (i.e., output embeddings, or LM Head) must be in full precision and with gradients activated, to get stable gradients into the residual stream. This introduces a non-negligible overhead, an issue avoided by ABFT as it does not supervise the final output. Moreover, fine-tuning attention projections without selectivity may cause harmful side effects toward the ICL out of the fine-tuned domain. We will compare the performance and efficiency of ABFT against end-to-end LoRA in the following experiments (Table 2) to highlight the efficiency and harmlessness of the ABFT method.
4 Main Experiments
In this section, we mainly confirm the effectiveness of the proposed ABFT, and find that: ABFT effectively improves the ICL performance to about 10%20% relatively, which requires the minimum parameters less than 0.05% to be full precision and gradient, with other parameters free to be quantized and gradient-free, and utilize 0.01% data cost compared to the previous works.
4.1 Experiment Settings
Models and Datasets.
We conduct our experiment on 9 modern LLMs: GPT2 (Large, XL) Radford et al. (2019), Falcon3 7B Team (2024b), Llama3 (8B, 43B, 56B) AI@Meta (2024), DeepSeek-R1 Distill Qwen 14B DeepSeek-AI (2025), Qwen2.5 32B Team (2024a); Yang et al. (2024), and SimpleScaling s1.1 32B Muennighoff et al. (2025); and 8 datasets: SST2, SST5 Socher et al. (2013), MR Pang and Lee (2005), Financial Phrasebank Malo et al. (2014), TREC Li and Roth (2002); Hovy et al. (2001), Subjective Wang and Manning (2012), Tweet Eval Emotion Mohammad et al. (2018), Tweet Eval Hate Basile et al. (2019) (Refer Appendix A.1 for details).
Hyperparameters.
We set: training samples , the number of demonstrations per ICL sample . A standard Adam optimizer Kingma and Ba (2014) is used with learning rate and pseudo-batch-size (i.e., we average gradients per samples before performing a single gradient step). We set the initial values , , and dynamically balance them with the PID algorithm (refer to Appendix A.3), stabilizing the number of attention heads identified as induction heads (see §5.3). The models are trained for steps.
Quantization Settings.
Models over 10B are quantized to 4-bit, with full-precision LoRA Hu et al. (2022) (inner dimension ) trained on and with learning rate .
Baselines.
Others.
We conduct all the experiments on a single NVIDIA A40 with 48GB VRAM. We repeat each experiment 4 times (10B) or 2 times (10B), and report the averaged results on 1024 fixed test inputs for each dataset. ICL-style inputs are built with library StaICC Cho and Inoue (2025).
4.2 Main Results
Model | Method | Param.* | Time | Acc | Acc |
Llama3 8.03B 4-bit, LoRA | Vanilla | - | - | 66.02 | |
E2E FT | 0.5B | 78.33 | 61.74 | ||
ABFT | 6.8M | 72.54 | 64.34 | ||
DeepSeek-R1 14.8B 4-bit, LoRA | Vanilla | - | - | 72.61 | |
E2E FT | 0.8B | 78.26 | 63.62 | ||
ABFT | 12M | 78.21 | 67.21 | ||
Qwen2.5 32.8B 4-bit, LoRA | Vanilla | - | - | 75.29 | |
E2E FT | 0.8B | 82.09 | 62.24 | ||
ABFT | 17M | 79.64 | 64.96 | ||
Llama3 55.6B 4-bit, LoRA | Vanilla | - | - | 70.95 | |
E2E FT | 1.1B | 82.80 | 64.86 | ||
ABFT | 33M | 79.52 | 67.32 |
Model | Template Consist. | Demonstration Consist. | ||
w/o ABFT | w/ ABFT | w/o ABFT | w/ ABFT | |
GPT2-XL | 81.28 | 91.74 | 68.38 | 82.75 |
Llama3 8B | 86.93 | 90.32 | 76.99 | 92.00 |
DeepSeek-R1 14B | 89.64 | 92.79 | 81.30 | 85.97 |
Qwen2.5 32B | 88.97 | 92.78 | 84.52 | 87.94 |
Llama3 56B | 92.78 | 93.90 | 82.10 | 87.49 |
The test results are shown in Table 1, where ABFT consistently outperforms all the baselines, even with enormous training sets (to MetaICL, ) and full-model fine-tuning (remind that ABFT only focuses on the and matrices), suggesting that ABFT is satisfyingly efficient in both time and data cost. Such results also provide strong empirical evidence for the effectiveness of induction heads in LLMs.
Towards Mechanistic Controllability.
To the best of our knowledge, ABFT is the first approach to train models without accessing final outputs, enabling model controlling via intermediate features or activations. Through this practice, we prototypically implement one of the visions of mechanistic interpretability Rai et al. (2024): by attributing the model’s inference to specific modules (circuits), we enable their local optimization, thereby improving overall performance effectively and efficiently.
Prediction Consistency.
We evaluate the prediction consistency against variations in (1) prompt templates and (2) demonstration sampling on StaICC-Diag Cho and Inoue (2025). For each query, we repeat predictions across different prompt templates and sampling strategies, and measure consistency as the ratio of the maximum consistent predictions (e.g., 6 positive vs. 3 negative yields ), averaged over the dataset; see Cho and Inoue (2025) and Appendix A.5 for implementation details. As shown in Table 3, ABFT significantly improves consistency across all 8 datasets, stabilizing ICL under diverse contexts and enhancing prompt design efficiency.
Prediction Bias against Wrong Labels.
Moreover, a known concern in ICL is the bias toward seen labels when ground-truth labels are absent in demonstrations (i.e., ) Zhao et al. (2021); Cho et al. (2025a), which can lead to incorrect predictions. Our testing on such scenario with and without ABFT in Fig. LABEL:fig:unseen shows that (see Appendix A.4 for experiment details): ABFT mitigates this issue via the , which penalizes incorrect labels during training and reduces the bias effects of induction heads. Notably, ABFT outperforms the 0-shot setting under unseen-label, suggesting the existence or emergence of unknown mechanisms that enable demonstrations in other categories to enhance ICL333Since that: such a phenomenon contrasts with existing views Cho et al. (2025a), where the explicit copying by the induction head is the only channel through which information is transferred from the demonstrations to the query (in Fig. 10, we show that the induction heads in ABFT model are almost fully suppressed in unseen-label scenario), where unseen-label demonstrations are harmful to ICL..
4.3 Comparison against End-to-end Fine-tuning
As mentioned before, end-to-end (E2E) fine-tuning on the in-domain dataset (not a wide dataset like MetaICL) with LoRA is also an alternative solution. However, in this section, we will show that compared with E2E fine-tuning, ABFT is more efficient, and more harmless on tasks out of the fine-tuning domain.
Time and Memory Cost.
Notice that in the E2E scenario, since the gradient is propagated from the final output of the model, the LM head, which is the top of the model, should be in full precision, to ensure a sufficient numerical resolution to utilize the mini-batch for mitigating the gradient noise in the stochastic gradient descent Hubara et al. (2016). This introduces a non-negligible overhead, as measured in Table 2. E2E fine-tuning slightly outperforms ABFT in in-domain accuracy (ACC), but incurs substantial training time and memory costs.
Harmlessness.
We evaluate out-of-domain (OD) performance on datasets different from the fine-tuned one (Table 2, ACC). Both ABFT and E2E fine-tuning degrade OD performance, but ABFT causes less harm. This supports a conclusion from Cho et al. (2025a): some induction heads are intrinsic and task-independent, while others are task-induced. ABFT on intrinsic heads harms OD performance, whereas ABFT on task-induced heads does not. In contrast, E2E fine-tuning on all heads broadly degrades OD performance.
Data Efficiency.
We test the accuracy against the training set size as a metric of data efficiency, for both ABFT and E2E fine-tuning, as shown in Fig. LABEL:fig:data_eff (refer to Appendix C.1 for results on other models). In the results, ABFT and E2E fine-tuning consistently benefit from more data samples, and in few-shot scenarios (), ABFT and E2E fine-tuning act equally, while E2E fine-tuning acts better when more training data is given. However, given the low-resource objective of ICL, and also the far more expensive time and memory cost of E2E fine-tuning, we can claim that ABFT has an advantage in the few-shot and low-resource scenario.
5 Analysis
5.1 Attention Visualization after ABFT
As shown in Fig. 4, we average the global (to ) and correct (to ) induction attention scores on the last token among attention heads and input samples on each transformer layer, on the validation set. Also, we provide a direct visualization of attention scores in Appendix C.3. Compared to the pre-trained model, the ABFT model tends to eliminate attention scores towards incorrect label tokens (), and shift the attention scores from the attention sinks (the first token) Xiao et al. (2024) and plain tokens to the correct label token, causing an enhancement to induction attention scores, continuously on the middle-to-last layers. Such visualization indicates that ABFT successfully generalizes to correct the behavior of attention heads.
5.2 Ablation Analysis
In this section, we disable some components utilized in the ABFT training protocol to suggest their necessity. The main results of such ablation experiments are shown in Table 4, where:
Model | Method | Time | Acc. |
Falcon3 7.46B | Vanilla | - | 67.55 |
ABFT | 79.59 | ||
w/o PID, | 76.86 | ||
w/o PID, w/o Head Filter | 75.57 | ||
w/o PID, | 72.13 | ||
w/o PID, | 56.47 | ||
Llama3 8.03B | Vanilla | - | 66.60 |
ABFT | 80.20 | ||
w/o PID, | 80.07 | ||
w/o PID, w/o Head Filter | 70.39 | ||
w/o PID, | 63.54 | ||
w/o PID, | 58.79 | ||
DeepSeek-R1 14.8B Dist. Qwen 4-bit, LoRA | Vanilla | - | 72.61 |
ABFT | 78.21 | ||
w/o PID, | 73.35 | ||
w/o PID, w/o Head Filter | 73.51 | ||
w/o PID, | 72.71 | ||
w/o PID, | 73.36 |
ABFT should be Localized.
In Table 4, disabling the head filter ((D) in Fig. 1) harms the accuracy. Knowing that all attention heads are trained to be induction heads under unfiltered ABFT loss, we can infer that: in LLMs, some attention heads with functions other than the induction head are still necessary for ICL, aligning with and enhancing the previous work Cho et al. (2025a). However, considering that some implicit antagonistic effects induced by unfiltered ABFT loss still promote the formation of other essential heads (i.e., when the ABFT loss from deeper heads propagates to shallower heads, its function becomes antagonistic with the ABFT loss directly connected to those shallow heads), the accuracy degradation with no head filters is not so significant.
Loss Factor ( and ) should be Balanced.
As mentioned in §4.1, we use the PID algorithm to adaptively balance the value of and in the loss calculation. In Table 4, we disable such adjustment and observe an accuracy drop. Especially, when we set the or to , the accuracy significantly degrades. This eliminates the doubt of “whether the loss factors are redundant” raised in §3, and we will discuss this in-depth in the next section (§5.3).
5.3 Balance the Loss Factor and
and Influence Induction Heads Antagonisticly.
As shown in Table 4, removing either loss term in Eq. 1 degrades performance, indicating that both are essential. Interestingly, the two terms introduce antagonistic implicit biases: the disperses attention across labels, reducing induction heads, while the concentrates attention on specific labels, increasing induction heads through a stepwise positive feedback loop. We track the number of induction heads during training on Llama3 8B and SST2 (see Appendix C.2 for more cases), as shown in Fig. 5, to support this observation.
Automatically Stabilizing Induction Head Number.
Ablation studies reveal that too many induction heads hinder fine-tuning and overall performance, as other functional heads are also needed for ICL; whereas an insufficient number prevents the model from handling ICL tasks. To maintain a stable number, we automatically adjust the antagonistic factor in Eq. 1 using a classical PID controller (Appendix A.3)444Since and are antagonistic, controlling alone suffices to stabilize induction head numbers.. As shown in Fig. 5, PID stabilizes induction head count, improves ABFT performance (Table 4), and reduces sensitivity to hyperparameters and , with accuracy remaining stable across settings (Fig. 6).
6 Consistency of Training Objective:
ABFT and End-to-end Fine-tuning
To explore whether the emergence of induction head is from ICL-style data—a key question in interpretability Chan et al. (2022); Reddy (2024); Singh et al. (2024)—we examine the consistency between ABFT and E2E objectives.
Principle.
Due to the lack of qualitative thresholds, it is hard to utilize statistic-based similarity measures to determine whether two models exhibit comparable similarity sufficient to indicate consistent training objectives. Therefore, our experiment is based on such a principle: if the fine-tuning terminations on both training objectives fall into the same basin of the loss function, then both fine-tuning trajectories are similar Neyshabur et al. (2020), so that the two training objectives are consistent. For such an end, we investigate the linear-connectivity Neyshabur et al. (2020); Ilharco et al. (2023) among the pre-trained parameters (as the start point of the fine-tuning) and fine-tuned parameters for E2E fine-tuning, and for ABFT. In detail, we mix these three parameters into a new model parameter set in the following form:
and then test the accuracy of as an anti-metric of model loss. As shown in Fig. 7 (upper), if the accuracy of mixed is better (or at least, not significantly worse) than the accuracy of and , we can infer that the and are in the same loss basin, with linear low-loss path observed.
Experiment and Result.
We conduct the aforementioned experiment protocol on SST2 and GPT2-XL. The results are shown in Fig. 7 (lower), showing no high-loss paths between the E2E fine-tuned model and the ABFT model. It suggests that they are in the same basin of the loss landscape, indicating the high similarity between these two training objectives. Moreover, we visualize the attention scores on ABFT and E2E fine-tuned models as shown in Fig. 8: compared with the pre-trained model, attention scores of the fine-tuned models on both objectives consistently focus on the correct label tokens, suggesting that E2E objectives imply a promotion to correct induction head. Moreover, as shown in Fig. 9, repeating the experiment on MetaICL shows MetaICL model lies in the same basin as ABFT, suggesting that full-model tuning on large datasets is essentially equivalent to ABFT, which can be seen as essential ICL fine-tuning.
7 Discussion
Conclusion.
In this paper, we propose a fine-tuning objective that strengthens the correctness of the induction head by accessing only the attention matrix, and demonstrate that it significantly improves the performance of ICL. Our results reinforce the induction head hypothesis for ICL interpretability and represent a first step toward controlling model behavior through mechanistic interpretability.
Towards Mechanistic Controllability.
In this paper, we raise the possibility of controlling the model’s behavior by some specific modules (often called circuit in the context of mechanistic interpretability), which opens up a new neural network model behavior-controlling paradigm: controlling only the modules that make significant contributions to the output, thereby substantially reducing the number of parameters that need to be adjusted and achieving excellent efficiency.
Limitations
Towards Open-end Tasks.
The first limitation lies in the fact that the induction head-based explanation of ICL Elhage et al. (2021); Singh et al. (2024); Cho et al. (2025a), so that our proposed ABFT approach, applies only to classification tasks with a finite label set. As mentioned in §4.2 and Fig. LABEL:fig:unseen, since our training objective consists of two factors, our approach is not limited to the simple retrieval setting where the ground-truth label appears in the demonstrations, while extending these methods to open-ended tasks remains an open challenge that requires further investigation on the basic mechanism. Nevertheless, given the current state of research on ICL interpretability, we have made full use of these findings and provided a valuable foundation for advancing model control through the scope of interpretability, i.e., Mechanistic Controllability.
Towards Better Mechanistic Controllability.
For our vision of Mechanistic Controllability, even though this paper successfully identifies a small set of modules (i.e., circuits) that require controlling towards better ICL performance, the control methods based on gradients and moderate amounts of data remain coarse. Therefore, future work could focus on gradient-free and data-free model editing, which directly edits some parameters utilizing a deeper understanding of the functional roles of model parameters.
Towards Better Performance.
It can be considered that some hyperparameters (see §4.1 and Appendix A.3), and the induction head filter (see §3) may be not optimal, restricting the performance. Discussing them in detail, and automatically optimizing them can be helpful for better performance of ABFT. Also, in Fig. 7, we observe that the extended line from the pre-trained model towards the ABFT model leads to better accuracy, suggesting a possibility of utilize model parameter to further improve accuracy without any gradient-based cost.
Towards Further Efficiency.
As shown in Appendix B, the and projections are significantly modified after ABFT only in some layers, that is, it is possible to further restrict the gradient-on parameters to some Transformer layers for better efficiency (notice that currently we activate the gradients of the attention mappings of all layers).
Acknowledgements
This work was supported by JST FOREST Program (Grant Number JPMJFR232K, Japan) and the Nakajima Foundation.
Appendix A Detailed Experiment Implementation
A.1 Model & Dataset Details
Dataset | Prompt Template (Unit) | Label Tokens |
SST2 | sentence: [input sentence] sentiment: [label token] n | negative, positive |
MR | review: [input sentence] sentiment: [label token] n | negative, positive |
FP | sentence: [input sentence] sentiment: [label token] n | negative, neutral, positive |
SST5 | sentence: [input sentence] sentiment: [label token] n | poor, bad, neutral, good, great |
TREC | question: [input sentence] target: [label token] n | short, entity, description, person, location, number |
SUBJ | review: [input sentence] subjectiveness: [label token] n | objective, subjective |
TEE | tweet: [input sentence] emotion: [label token] n | anger, joy, positive, sad |
TEH | tweet: [input sentence] hate speech: [label token] n | normal, hate |
Model | Repository |
GPT2-L | openai-community/gpt2-large |
GPT2-XL | openai-community/gpt2-xl |
Falcon3 | tiiuae/Falcon3-7B-Base |
Llama3 (8B) | meta-llama/Meta-Llama-3-8B |
DeepSeek-R1 | deepseek-ai/DeepSeek-R1-Distill-Qwen-14B |
Qwen2.5 | Qwen/Qwen2.5-32B |
SimpleScaling s1.1 | simplescaling/s1.1-32B |
Llama3 (43B) | chargoddard/llama3-42b-v0 |
Llama3 (56B) | nyunai/nyun-c2-llama3-56B |
Models.
All the models in this paper are loaded from huggingface. In detail, we list the huggingface repository name to keep the repeatability of this paper, as shown in Table 6.
Dataset Split.
As also described by Cho and Inoue (2025), we randomly sample 1024 data samples from the original dataset to build the inputs for training, and sample (especially, for FP dataset, for TEH dataset) data samples for the demonstrationsqueries for the testing, respectively.
Demonstration Sampling.
To generate the training examples of demonstrations, we randomly sample data examples from the aforementioned 1024 data, and concatenate them into the inputs, with the prompt templates shown in §A.2. To generate the testing examples, for each query in the 512 samples, we sample two sequences of demonstrations from the 4096 data samples, and concatenate them into testing inputs, 2 for one query sample, so that 1024 for one dataset.
A.2 Prompt Templates
A.3 Details of PID Algorithm
On each model update step (i.e., when the gradients from all the samples of the -th pseudo batch (of data samples) are propagated), we calculated the identified induction head numbers from the induction head filter described in §3 and Fig. 1 averaged on the data samples as . Given the similar averaged induction head number on the previous time step as , we can calculate the updated term555Remind that for the sake of simplicity, we only control , given the findings of and are antagonistic, as shown in §5.3. by standard PID algorithm as:
where the , , and are hyperparameters. By such calculation, we implement a feedback control to stabilize the number of induction heads among training steps.
A.4 Experiment Protocol of Unseen Label
In Fig. LABEL:fig:unseen, we examine that ABFT model can utilize the demonstration with wrong label to improve the ICL performance. Here we introduce the experiment protocol.
First, we train a Llama3 8B on TREC (a 6-way classification dataset) with the ABFT method. Then, to test the trained ABFT model on the unseen label condition, we build special test inputs: for each query with label , we choose demonstrations with label , and utilize the standard template shown in Table 5 to build the inputs, then test the accuracy. Notice that during the training, no special sampling for the inputs is conducted, i.e., the training is not under the unseen label setting, so that such an experiment protocol also confirms the generalization of ABFT methods on a different distribution. Moreover, we repeat the induction head visualization shown in Fig. 4 on the unseen label condition in Fig. 10, where the induction heads in the ABFT model are almost fully suppressed but with considerable inference accuracy, which implies a new inference mechanism.
A.5 Experiment Protocol of Stability against Prompting
In Table 3, we test whether the prediction of ICL is stable against various (1) prompt templates and (2) demonstration sampling, on the StaICC-Diag Cho and Inoue (2025) benchmark, whose method is described briefly below.
Method.
To test the prediction robustness against prompt templates / demonstration sampling, we repeat several predictions for each query on various prompt templates / demonstration sampling, and calculate the ratio of the maximum consistent group (e.g., we get 6 positive and 3 negative predictions on one query, then the ratio is ). The robustness metrics are the average value of the whole dataset. Refer to Cho and Inoue (2025) for the detailed implementation. Notice that only the consistency is tested in these experiments, without observing the accuracy.
Result.
The robustness metrics among prompt templates / demonstration sampling averaged on all 8 datasets before and after ABFT are shown in Table 3, where both terms of the robustness are significantly improved after ABFT, suggesting that ABFT stabilizes ICL for various contexts, providing higher efficiency on prompt designing. Also, given the results with mitigating prediction sensitivity and bias against prompt templates / demonstration sampling, which is consistent with the objective of output calibration Zhao et al. (2021); Fei et al. (2023); Han et al. (2023); Zhou et al. (2024); Jiang et al. (2023); Cho et al. (2025b), ABFT can be regarded as an implicit calibration inside the LLM.
Appendix B Parameter Shift after ABFT against Layers
We utilize the Frobenius norm to visualize the shifting distance of the parameter matrix before and after ABFT () as . The results are shown in Fig. 17, 18, 19, where, although each model exhibits its own pattern in terms of distance across layer numbers, certain layers consistently show significantly lower distances within every model.
Moreover, even though the early layers accumulate more gradients (since the gradients from each later layer propagate backward to them), the peak of the shifting distance typically appears in the middle to later layers. This observation is consistent with previous works on Induction Heads Cho et al. (2025a).
Appendix C Augmentation Results
C.1 Augmentation Results for Data Efficiency (Fig. LABEL:fig:data_eff)
We repeat the data efficiency experiments shown in Fig. LABEL:fig:data_eff on Qwen2.5 32B, as shown in Fig. 15. The results are globally consistent with Fig. LABEL:fig:data_eff.
C.2 Augmentation Results for Number of Induction Heads against Training Processing (Fig. 5)
C.3 Augmentation Results for Attention Visualization (Fig. 4 and 8)
As shown in Fig. 16, we visualize the attention score on the last token of the given input example in the validation set on Llama3 8B, and repeat this visualization on more input cases for Fig. 16 and Fig. 8 in Fig. 29. Moreover, we expand the Fig. 16 towards more layers in Fig. 30, and Fig. 8 in Fig. 31. We observe that ABFT significantly modifies the attention distribution in the middle layers, while in the early and late layers, neither ABFT nor E2E has a substantial impact on attention scores.
Appendix D Statements
Author Contributions Statement.
Hakaze Cho, also known as Yufeng Zhao, handled the entire workload in this paper. He provided ideas, designed / conducted experiments, collected / described data, and wrote / revised the paper.
Peng Luo suggested the PID algorithm, which is an essential part of this paper, and provided some technical support for it.
Naoya Inoue, as the supervisor, provided valuable feedback, revisions, and financial support for this paper.
M.K. and R.K. made minor contributions that placed them on the borderline of authorship criteria: they participated in discussions and offered comments, some of which were incorporated.
License for Artifacts.
Models and datasets used in this paper are used in their original usage, and are open-sourced with the following license:
mit: GPT2-L, GPT2-XL, DeepSeek-R1
Individual License or Unknown: Falcon3, Llama3 8B, Llama3 43B, Llama3 56B, SST5, MR, TREC, SUBJ, TEE, TEH
Apache 2.0: Qwen2.5, SimpleScaling s1.1
cc-by-sa-3.0: SST2, FP,
AI Agent Usage.
AI Agents are used and only used for writing improvement in this paper.